You are investigating a rare genetic brain disorder and would like to identify the gene responsible, clone it and determine its function. A very small protein isolated from the brains of healthy individuals who have the disease. The protein has a molecular weight of about 2700 Daltons. Assume that the molecular weight of a typical amino acid is 100 Daltons.

  1. How many amino acids do you expect the protein to be?
  2. How long, in nucleotides, would the corresponding mature mRNA be from the beginning of the start codon to the end of the stop codon?
  3. Mature mRNA already has introns spliced out. Look back to your notes from Bio 204 – in your own words, what is an intron? What is an exon?
  4. You find that your mature mRNA is longer than the length you indicated in part B. What most likely explains this difference in length? Explain your reasoning.
Answer #1

a) molecular weight of the protein is 2700 Da, and molecular weight of an amino acid is 100 Daltons so the number of amino acids= mass of protein/mass of an amino acid=2700/100=27

b) one codon codes for amino acid and one codon is 3 nucleotides long, so the number of nucleotides that code for amino acids= 327=81

there is one stop codon at the end so the total number of nucleotides from the start codon to stop codon= 81+3=84

c) introns are sequences found in the pre-mRNA which are spliced out, it is not part of the mature mRNA.

exons are found in between the intron sequences in the pre-mRNA, introns are spliced out and exons are joined together to form the mature mRNA.

d) the mature mRNA has 5`-UTR and 3`-UTR, 5`-UTR is the untranslated region in the 5` end and 3`-UTR is untranslated region in the 3` end , which included poly A tail, so the mature mRNA is longer than the open reading frame ( ORF, nucleotides from start codon to stop codon).